OpenStudy (anonymous):

Solve the system algebraically. Check your work. 5x + 2y = 10 3x + 2y = 6

OpenStudy (anonymous):

This answer is x=2, y=0

OpenStudy (anonymous):

thanks

OpenStudy (tkhunny):

@clayton_reed Prove it. Don't buy it!

OpenStudy (anonymous):

\[(5x+2y) - (3x+2y) = 2x\] \[(5x+2y) - (3x+2y) = 10 -6 = 4\] Therefore \[2x = 4\] x=2 \[5\times 2 + 2y = 10\] \[2y = 10-10 = 0\] y=0

OpenStudy (anonymous):

Can you understand it, tkhunny?

OpenStudy (tkhunny):

I don't understand why clayton_reed gets the answer and is not required to show ANY of his own work. Other than that, it seems okay.